# Supporting

### v1

@@ 5 Real data analysis
SCL-Linear $285.2(12.0)$ $\boldsymbol{384.8}(12.6)$ $501.3(12.2)$
SCL-Gaussian $\boldsymbol{285.4}(11.4)$ $\boldsymbol{384.8}(11.2)$ $\boldsymbol{502.1}(12.2)$
Wang’s method $276.1(12.2)$ $366.6(14.8)$ $476.7(14.4)$
QIQ-learning $284.1(12.6)$ $379.6(13.9)$ $493.8(11.8)$
6 Quantile optimal dynamic treatment regimes
All previous discussions focus on single-stage quantile OTRs. In managing chronic diseases, it is often necessary to adapt treatment strategies over time based on a patient’s response to prior interventions. This motivates the estimation of quantile optimal dynamic treatment regimes (ODTRs) using data from longitudinal studies.
For brevity, consider a two-stage longitudinal study. Let $A_{1}\in\{0,1\}$ denote the treatment at stage $1$ and $A_{2}\in\{0,1\}$ denote the treatment at stage 2. The baseline covariates are denoted by $\boldsymbol{X}_{1}$ at stage $1$ and $\boldsymbol{X}_{2}$ at stage $2$. We are interested in improving the continuous outcome $Y$ at the end of the study. Let $\boldsymbol{H}_{1}=(\boldsymbol{X}_{1})$ and $\boldsymbol{H}_{2}=(\boldsymbol{X}_{1},A_{1},\boldsymbol{X}_{2})$ be the history information at stage $1$ and $2$, respectively. A dynamic treatment regime (DTR) is $d=(d_{1},d_{2})$, where $d_{k}$ can depend on $\boldsymbol{H}_{k}$, $k=1,2$. To study the ODTRs, we utilize the counterfactual outcome framework. Suppose $Y(a_{1},a_{2})$ is the counterfactual outcome that the subject receives treatment $(A_{1},A_{2})=(a_{1},a_{2})$, $a_{1},a_{2}\in\{0,1\}$, and $\boldsymbol{X}_{2}(a_{1})$ is the counterfactual covariates that the patient receives treatment $A_{1}=a_{1}$, $a_{1}\in\{0,1\}$. Given the DTR $d$, the corresponding counterfactual history at stage $2$ is $\boldsymbol{H}_{2}(d_{1})=[\boldsymbol{X}_{1},d_{1}(\boldsymbol{H}_{1}),\sum_{a_{1}}\boldsymbol{X}_{2}(a_{1})I\{d_{1}(\boldsymbol{H}_{1})=a_{1}\}]$, and the corresponding counterfactual outcome is $Y(d)=\sum_{a_{1},a_{2}}Y(a_{1},a_{2})I[d_{1}(\boldsymbol{H}_{1})=a_{1},d_{2}\{\boldsymbol{H}_{2}(d_{1})\}=a_{2}]$. With a slight abuse of notation, we continue to use $d^{*}$ to define the quantile ODTRs, i.e., $d^{*}=\operatorname*{arg\,max}_{d}Q_{\tau}\{Y(d)\}$. We adopt the following assumption Wang et al., (2018).
Assumption 1: (i) $Y=\sum_{a_{1},a_{2}}Y(a_{1},a_{2})I(A_{1}=a_{1},A_{2}=a_{2})$ and $\boldsymbol{X}_{2}=\sum_{a_{1}}\boldsymbol{X}_{2}(a_{1})I\{A_{1}=a_{1}\}$; (ii) $Y(a_{1},a_{2}),\boldsymbol{X}_{2}(a_{1})\perp\!\!\!\perp A_{1}\mid\boldsymbol{H}_{1}$ and $Y(a_{1},a_{2})\perp\!\!\!\perp A_{2}\mid\boldsymbol{H}_{2}$ for $a_{1},a_{2}\in\{0,1\}$; (iii) $10<P(A_{k}=a_{k}|\boldsymbol{H}_{k})<$ 1 for $k\in\{1,2\},a_{k}\in\{0,1\}$.
We first consider the identification of the quantile ODTRs. Since $Q_{\tau}\{Y(d)\}=\sup\{q:S(q,d)>1-\tau\}$, it suffices to identify the survival function $S(q,d)=P\{Y(d)>q\}$ for any $q$ and $d$ in order to identify the quantile ODTRs. Define $\pi^{*}=(\pi_{1}^{*},\pi_{2}^{*})$, where $\pi_{k}^{*}(a|\boldsymbol{h}_{k})=P(A_{k}=a_{k}|\boldsymbol{H}_{k}=\boldsymbol{h}_{k})$ is the propensity score at stage $k$, $k=1,2$. Under Assumption 1, the $S(q,d)$ can be identified as Zhao et al., 2015a
${S}(q,d)=E\left[\frac{I(Y>q)I\{d_{1}(\boldsymbol{H}_{1})=A_{1}\}I\{d_{2}(\boldsymbol{H}_{2})=A_{2}\}}{\pi_{1}^{*}(A_{1}|\boldsymbol{H}_{1})\pi_{2}^{*}(A_{2}|\boldsymbol{H}_{2})}\right].$ (9)
For brevity, we present the identification result for $S(q,d)$ based solely on the propensity score, and the estimator for the quantile ODTRs below is also based on inverse propensity weighting. While it is feasible to construct a doubly robust estimator for the quantile ODTRs within our sequential classification learning framework, doing so would involve considerable complexity due to the need for a backward iterative procedure Zhang and Zhang, (2018). Therefore, to convey the core idea of our method in the longitudinal setting, we focus on the estimator based on the propensity score model.
We now discuss the estimation of the quantile ODTRs using Algorithm 1. It suffices to provide estimators of $d_{q}^{*}=\operatorname*{arg\,max}_{d}S(q,d)$ and $\max_{d}S(q,d)=S(q,d_{q}^{*})$. We first focus on the estimation of $d_{q}^{*}$. Suppose the observed data is $\{\boldsymbol{X}_{i1},\boldsymbol{X}_{i2},Y_{i},A_{i1},A_{i2}\},i=1,...,n$, and assume $\pi^{*}$ is known. Note that the regimes $d_{1}(\boldsymbol{H}_{1})$ and $d_{2}(\boldsymbol{H}_{2})$ can always be represented as $I\{f_{1}(\boldsymbol{H}_{1})>0\}$ and $I\{f_{2}(\boldsymbol{H}_{2})>0\}$ for some decision functions $f_{1},f_{2}$. Therefore, utilizing concave surrogate loss to replace the 0-1 loss $I\{d_{1}(\boldsymbol{H}_{1})=A_{1}\}I\{d_{2}(\boldsymbol{H}_{2})=A_{2}\}$ in the identification (9) Zhao et al., 2015a , we can estimate the ODTRs $d_{q}^{*}$ via $\hat{d}_{q}=I(\hat{f}_{q}>0)=\{I(\hat{f}_{1q}>0),I(\hat{f}_{2q}>0)\}$, where
$\hat{f}_{q}=(\hat{f}_{1q},\hat{f}_{2q})=\operatorname*{arg\,max}_{f_{1},f_{2}\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\frac{I(Y_{i}>q)\varphi\{(2A_{i1}-1)f_{1}(\boldsymbol{H}_{i1}),(2A_{i2}-1)f_{2}(\boldsymbol{H}_{i2})\}}{\pi_{1}^{*}(A_{i1}|\boldsymbol{H}_{i1})\pi_{2}^{*}(A_{i2}|\boldsymbol{H}_{i2})}$ (10)
$-\lambda_{n}(\|f_{1}\|^{2}+\|f_{2}\|^{2}).$
Here, $\varphi(x,z)=\min\{x-1,z-1,0\}+1$ for $x,z\in\mathbb{R}$ is the concave surrogate loss, and the penalty term $\lambda_{n}(\|f_{1}\|^{2}+\|f_{2}\|^{2})$ is introduced to prevent overfitting. The function space $\mathcal{F}$ and the norms $\|f_{1}\|^{2},\|f_{2}\|^{2}$ can be chosen as that in Section 2.2. For a discussion on solving the maximization problem (10), we refer to Zhao et al., 2015a .
Next, we consider the estimation of $S(q,d_{q}^{*})$. Given that $\hat{d}_{q}=I(\hat{f}_{q}>0)$ is an estimator of $d_{q}^{*}$, we can estimate $S(q,d_{q}^{*})$ as follows:
$\widehat{S}_{h_{n}}(q,\hat{f}_{q})=\frac{1}{n}\sum_{i=1}^{n}\frac{I(Y_{i}>q)l[A_{i1},\Phi\{\hat{f}_{1}(\boldsymbol{H}_{i1})/h_{n}\}]l[A_{i2},\Phi\{\hat{f}_{2}(\boldsymbol{H}_{i2})/h_{n}\}]}{\pi_{1}^{*}(A_{i1}|\boldsymbol{H}_{i1})\pi_{2}^{*}(A_{i2}|\boldsymbol{H}_{i2})},$
where $\Phi(x),x\in\mathbb{R},$ is the cumulative distribution function of the standard normal distribution, and $l(x,z)=xz+(1-x)(1-z)$ for $x,z\in\mathbb{R}$. We can then apply Algorithm 1 to estimate the quantile ODTRs by using the $\hat{f}_{q}$ and $\widehat{S}_{h_{n}}(q,\hat{f}_{q})$ as defined in this section.
@@ 7 Discussion
In this paper, we introduce a sequential classification learning method for estimating quantile OTRs. Our approach leverages the benefits of classification learning, both in optimization and statistical properties, to overcome the challenges posed by non-convexity and regime linearity inherent in previous methods. To tackle the inconsistency and ineffectiveness issues associated with discrete outcomes, we propose to smooth the quantile. Our method demonstrates satisfactory performance in both simulation and real data.
This paper focuses on a binary treatment scenario. In certain diseases, multiple treatment options may be available. It would be of interest to extend our method to multiple treatments. We can adopt the idea from Zhang et al., (2020) to embed all possible treatments into a series of points in Euclidean space and combine the angle-based method with Algorithm 1 to learn the quantile OTRs. The angle-based method is similar to classification learning, which utilizes machine learning techniques to enhance computational efficiency and flexibly accommodate nonlinear regimes.
In various applications, the baseline covariates may be large, and variable selection is important to have an efficient and explanatory estimation. We can incorporate the LASSO or SCAD penalty Tibshirani, (1996); Fan and Li, (2001) to choose variables for the linear quantile OTRs. If we aim to derive a sparse nonlinear regime with the Gaussian kernel, the gradient-based variable selection method Yang et al., (2016) can be employed.

### v2

@@ 5 Real data analysis
SCL-Linear $281.8(4.4)$ 380.2(8.5){380.2}(8.5) $496.4(11.3)$
SCL-Gaussian $\boldsymbol{283.3}(4.8)$ $\boldsymbol{383.2}(8.3)$ $\boldsymbol{500.2}(9.8)$
Wang’s method $276.9(5.9)$ $368.1(9.8)$ $478.3(10.8)$
QIQ-learning $280.3(4.4)$ $375.9(8.0)$ $490.6(8.3)$
To strengthen our finding, we also add another metric: the Rand index (RI) rand1971objective, which quantifies the stability of different methods in assigning quantile optimal treatments. Specifically, we randomly divide the data into two subsets, the training and test sets, allocating 80% of the observations to the former and the remaining 20% to the latter. Within the training set, we further partition the data into two equal halves and obtain two estimated quantile OTRs from each half, denoted by $\hat{d}_{1}$ and $\hat{d}_{2}$, respectively. We then compute the Rand index on the test set as $\sum_{j=1}^{m}I\{\hat{d}_{1}(\boldsymbol{X}_{j})=\hat{d}_{2}(\boldsymbol{X}_{j})\}/m$. The closer the Rand index is to 1, the greater the stability of the method. Repeating this process 200 times, we report the average Rand index for different methods in Table 3. It is evident that our method with a Gaussian kernel achieves the highest stability in assigning quantile optimal treatments, as reflected by a larger Rand index relative to competing approaches.
Table 3: Mean (standard deviation) of Rand index (“RI”) for different methods. The best result in each scenario is highlighted in bold.
RI for $\tau=0.25$ RI for $\tau=0.50$ RI for $\tau=0.75$
SCL-Linear $0.657(0.223)$ $0.823(0.202)$ $0.813(0.195)$
SCL-Gaussian $\boldsymbol{0.801}(0.162)$ $\boldsymbol{0.881}(0.151)$ $\boldsymbol{0.870}(0.178)$
Wang’s method $0.523(0.109)$ $0.554(0.108)$ $0.605(0.166)$
QIQ-learning 0.614(0.074){0.614}(0.074) $0.645(0.068)$ $0.643(0.073)$
@@ 6 Discussion
In this paper, we introduce a successive classification learning method for estimating quantile OTRs. Our approach leverages the benefits of classification learning, both in optimization and statistical properties, to overcome the challenges posed by non-convexity and regime linearity inherent in previous methods. To tackle the inconsistency and ineffectiveness issues associated with discrete outcomes, we propose to smooth the quantile. Our method demonstrates satisfactory performance in both simulation and real data. We further extend our method to longitudinal studies and survival data in Section S4 of the supplementary material to enhance the applicability of our framework.
This paper focuses on a binary treatment scenario. In certain diseases, multiple treatment options may be available. It would be of interest to extend our method to multiple treatments. We can adopt the idea from zhang2020multicategory to embed all possible treatments into a series of points in Euclidean space and combine the angle-based method with Algorithm 1 to learn the quantile OTRs. The angle-based method is similar to classification learning, which utilizes machine learning techniques to enhance computational efficiency and flexibly accommodate nonlinear regimes.
In various applications, the baseline covariates may be large, and variable selection is important to have an efficient and explanatory estimation. We can incorporate the LASSO or SCAD penalty tibshirani1996regression; fan2001variable to choose variables for the linear quantile OTRs. If we aim to derive a sparse nonlinear regime with the Gaussian kernel, the gradient-based variable selection method yang2016modelfree can be employed.